OpenStudy (anonymous):

Sarah has a total of 52 DVDs and CDs. If the number of CDs is two more than four times the number of DVDs, how many CDs does she have?

OpenStudy (anonymous):

let x = number of CDS let y = number of DVDS 4x+2+4y=52 4y+2+4y=52 8y=50 y=50/8 y=6.25 number of CDS= 4x + 2 = 4(6.25) + 2 = 27

OpenStudy (anonymous):

Wait so you get 4x= 25 and x is 6.25...? @JayDS

OpenStudy (anonymous):

What is that answer? @JayDS

OpenStudy (anonymous):

yes, that is correct.

OpenStudy (anonymous):

I was even able to double check the answer.

OpenStudy (anonymous):

So she has 6.25 CDs....? Haha

OpenStudy (anonymous):

how? by subbing the values back into the equation. Yeh I find it weird as well haha, but it works, let me show you.

OpenStudy (anonymous):

let x = number of CDS << 4x+2 we can make it 4y + 2, because we have to make it in terms of number of DVDS(y) to solve it. let y = number of DVDS <<4y 4y+2+4y=52 since we figured out that y=6.25 4(6.25) + 2 + 4(6.25) = 52

OpenStudy (anonymous):

Oh ok. I see. So she definitely has 6.25 CDs. The answer is 6.25 correct? Lol just making sure!! @JayDS

OpenStudy (anonymous):

well it works lol.

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

but not that it practically makes sense in reality unless you think of it this way that the 0.25 is 1/4 of a broken CD.